OpenStudy (anonymous):

If (x + 1)(x - 3) = 5, then which of the following statements is true?

OpenStudy (anonymous):

(x+1)(x-3)=5 =>x^2-3x+x-3=5 =>x^2-2x-3=5 =>x^2-2x-8=0 x^2-4x+2x-8=0 =>x(x-4)+2(x-4)=0 =>(x-4)(x+2)=0 x=4,-2